OpenStudy (anonymous):

how to find log .0078654 ?

OpenStudy (mathstudent55):

What are you allowed to use? A calculator, tables?

OpenStudy (anonymous):

Tables

OpenStudy (kirbykirby):

Oh my lol I never had to use log tables

OpenStudy (kirbykirby):

(Texbtooks still print these?)

OpenStudy (kirbykirby):

Textbooks*

OpenStudy (anonymous):

I can answer in calculator but when i'm using the tables. Argh, cant find the answer..

OpenStudy (anonymous):

Do you think changing it to Exponential Notation would help you out?

OpenStudy (mathstudent55):

Assuming base 10 log .0078654 = log( (7.8654) x (10^-3) )

OpenStudy (unklerhaukus):

\[ \log (0.0078654)=\log (7.8654\times10^{-3})\]\[\qquad\qquad=\log (7.8654)+\log(10^{-3})\]\[\qquad\quad=\log (7.8654)-3\log(10)\]\[\quad\quad=\log (7.8654)-3\]

OpenStudy (mathstudent55):

log .0078654 = log( (7.8654) x (10^-3) ) = log(7.8654) + log(10^-3) = log(7.8654) look up in table + (-3)